OpenStudy (anonymous):

If f(x)=log3 (x+1), what is f-1(2)

OpenStudy (anonymous):

3log(3)

OpenStudy (anonymous):

The answers are 1,8,10,27

OpenStudy (freckles):

You found f(2) @asdfghjkl24562

OpenStudy (freckles):

solve for x \[2=\log_3(x+1)\]

OpenStudy (anonymous):

ohhhhhhh

OpenStudy (michele_laino):

hint: using the definition of logarithm, we can write: \[\Large x + 1 = {3^y}\]

OpenStudy (carolinar7):

Oh the logarithmic is the in the inverse of exponential functions, so all you have to do is cinvert it to an exponential function and substitute the velue

OpenStudy (anonymous):

You guys are amazing

OpenStudy (carolinar7):

\[(2)+1=3^{y}\]

OpenStudy (michele_laino):

thanks!

OpenStudy (carolinar7):

3=3\[^{y}\]

OpenStudy (carolinar7):

\[3 = 3^{y}\]

OpenStudy (freckles):

it's suppose to be \[3^2=x+1\]

OpenStudy (carolinar7):

no

OpenStudy (freckles):

\[f^{-1}(2) =x \text{ implies } f(x)=2 \\ \text{ so we need to solve } 2=\log_3(x+1) \\ \text{ and converting to exponential form gives } \\ 3^2=x+1\]

OpenStudy (carolinar7):

Oh ya